A trust fund is planning to invest up to $6000 in two types of bonds.: A and B. Bond A is safer than bond B and carries a dividend of 8% while bond B carries a dividend of 10%. Suppose that the bond’s rule state that no more than $4000 may be invested in bond B, while at least $1500 must be invested in bond A. The amount invested in bond B cannot exceed one half the amount invested in bond A. How much should be invested in each type of bond to maximize the fund’s return?


1
Expert's answer
2021-03-16T08:38:53-0400

"x=" amount invested in A bonds

"y=" amount invested in B bonds


"x+y\leq6000"

"0\leq y\leq4000"

"x\geq1500"

"y\leq\dfrac{1}{2}x"

Our linear optimization problem is:

Maximize "z=0.08x+0.1y" subject to


"x+y\leq6000"

"0\leq y\leq4000"

"x\geq1500"

"y\leq\dfrac{1}{2}x"


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Segment & Equation & z=0.08x+0.1y \\ \hline
 AB & x=1500, & z=120+0.1y\\
 & 0\leq y\leq 750 & \\
 \hdashline
 BC & y=\dfrac{1}{2}x, & z=0.13x \\
& 1500\leq x\leq 4000 & \\
 \hdashline
 CD & y=6000-x, & z=600-0.02x \\
& 4000\leq x\leq 6000 & \\
 \hdashline
 DA & y=0, & z=0.08x \\
& 1500\leq x\leq 6000 & \\
 \hdashline
\end{array}"



"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Point & z=0.08x+0.1y \\ \hline
 A(1500, 0) & z=120\\

 \hdashline
 B(1500, 750) & z=195 \\

 \hdashline
 C(4000, 2000) & z=520 \\

 \hdashline
 D(6000, 0)& z=480 \\
 \hdashline
\end{array}"



Because the point "(4000, 2000)" produces the highest fund’s return we conclude that $4000 should be invested in


How much should be invested in bond A and $2000 should be invested in bond B.